one integer is 6 less than another integer. the product of the lesser integer and -7 is 42 more than the greater integer. what are the integers?
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement:
:
one integer is 6 less than another integer.
x = y - 6
:
the product of the lesser integer and -7 is 42 more than the greater integer.
-7*x = y + 42
:
what are the integers?
:
Substitute (y-6) for x in the above equation, find y:
-7(y-6) = y + 42
-7y + 42 = y + 42
Subtract 42 from both sides
-7y = y
y = 0
:
Find x
x = 0 -6
x = -6
:
:
Check solution in the statement:
the product of the lesser integer and -7 is 42 more than the greater integer.
-7(-6) = 0 + 42